OpenStudy (anonymous):

The line including the two points (5, 3), (8, -2)

OpenStudy (btaylor):

To find the equation, 1. Find the slope. m=(change in y)/(change in x)=(-2-3)/(8-5) = -5/3 2. Plug in one value into y=mx+b form and solve for b. 3 = 5(-5/3)+b 3 = -25/3+b 34/3=b so your equation is y=(-5/3)x+34/3

OpenStudy (btaylor):

Do you understand?

OpenStudy (anonymous):

yes but the answer does not match the one i have

OpenStudy (btaylor):

What do you have?

OpenStudy (anonymous):

can you help me with this one? The line with m = 2, intersecting 2x - 3y = 6 at x = 3.

OpenStudy (btaylor):

OK...we have a line y=2x+b. We know that, at x=3, y=0 (2x3 - 3x0 = 6). So, we can use the point (3,0) to find b. 0 = 2(3) + b --> b = -6 Your result: y = 2x - 6

OpenStudy (austinl):

Hey, @BTaylor for future reference, you should guide them to the answer. Not merely work it out for them and give them the answer.

OpenStudy (btaylor):

@austinL, that is "guiding her to the answer". I'm not writing "y = 2x -6" and then explaining it (if I do at all), but showing how to work it out.

OpenStudy (austinl):

Still, it is better to guide step by step.

OpenStudy (btaylor):

That's a nice opinion for you to have. I'll do what works for me.

OpenStudy (austinl):

You should do what complies with the Code Of Conduct as well as what is the accepted convention among the moderators.